QueenMidnight16:

i need help fast The work of a student to solve the equation 4(2x − 4) = 8 + 2x + 8 is shown below: Step 1: 4(2x − 4) = 8 + 2x + 8 Step 2: 6x − 8 = 16 + 2x Step 3: 6x − 2x = 16 + 8 Step 4: 4x = 24 Step 5: x = 6 In which step did the student first make an error and what is the correct step? Step 2; 8x − 4 = 2(6 + x + 6) Step 2; 8x − 16 = 16 + 2x Step 3; 6x − 2x = 16 − 8 Step 3; 6x + 2x = 16 + 8

Walen:

i feel like its step 2 because i dont understand how 4(2x-4) = 8 it should be 8x-16 = 8

Walen:

but it seems like u caught up to that

Walen:

with ur 6x + 2x = 16+8 u can do 8x = 27 x = 27/8 x=3

Walen:

now give me a medal and pet me on the head master ;)

Walen:

ask me if u dont understand anything